DX Actuator Testing

16:26
  • Loop Closed
  • Reconstructor 20130919_014135.fits
  • Gains 1.06 0.29 0.20
  • Optical Loop Data: 20170502_163337
  • Took Flat test_163337_400.sav
  • Re-applied it with 230 modes
  • Applied Disturb /local/aomeas/adsec_calib/CMD/disturb/old_KL_v16/dist_flao1_ts4_atm_s0.8_L040.0_v15.0_ovfreq1000.00_sd1983.fits
  • Gains: 1.01 0.49 0.59
  • Optical Loop: 20170502_164817, 164900
  • Gains back to 1.06 0.29 0.20
  • Optical Loop 20170502_165034
  • Saved flat: test_165034_400.sav
  • re-applied with 222 modes
17:13
  • put actuators 474, 487, 371 in no icc list.
  • That made things worse.
  • The above were accidentally in no pos
17:57
  • put 474, 487, and 371 into no icc list. (the real one)
  • Closed Loop 1.12 0.20 0.20
  • Dataset 181102
  • new flat test_181102_400.sav
  • Loaded with 245 modes.
18:13
  • Removed 474, 487, 371 from list
  • Take out 599 from "no_pos" list
  • Gain: 0.85 0.20 0.28
  • Dataset: 182718
  • Flat: test_1812718_400.sav (not a typo)
  • re-applied with 258 modes
18:31
  • Removed 562 from "no_pos" list
  • Loaded test_1812718_400.sav with 171 modes
  • Gain 0.85 0.28 0.20
  • Dataset: 183955
  • Flat: test_183955_400.sav
  • re-applied with 257 modes
18:46
  • Removed 559 from "no_pos" list
  • 120 modes on default flat
  • 216 with previous flat
  • Gain 0.86 0.20 0.20
  • Dataset: 185818
  • Flat: test_185818_400.sav
  • re-applied with 257 modes
19:04
  • Removed 558 from "no_pos" list
  • default flat 136 modes
  • last flat: 229 modes
  • Gains: 0.87 0.20 0.20
  • Dataset: 191102
  • Flat: test_191102_400.sav
  • re-applied with 274 modes
19:18
  • Removed 379 from "no_pos" list
  • Can't load proagrams after multiple tries, 274993 um reading.
19:28
  • Removed 378 from "no_pos" list
  • Can't load programs after multiple tries, 263615 um reading
19:42
  • Removed 376 from "no_pos" list
  • Default flat 139 modes
  • Last flat 174 modes
  • Gains: 0.84 0.20 0.20
  • Dataset: 195523
  • flat: test_195523_400.sav
  • re-applied with 276 modes
  • 270 modes after rest and reset.
20:04
  • Removed 560 and 561 from "no_pos" list
  • default flat 111 modes
  • Last flat 145 modes
  • Gains: 0.96 0.28 0.20
  • Dataset: 201502
  • Flat: test_201502_400.sav
  • re-applied 245 modes
  • New Flat: test_202257_400.sav
  • re-applied 230 modes
  • Will work on these some more.

-- %USERSIG{GregoryTaylor - 2017-05-02}%

Comments:

No Comment
Topic revision: r3 - 02 May 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback